OpenStudy (anonymous):

how do you find the solutions for y=x2+4x-5 and y=x2-2?

OpenStudy (amorfide):

they give you the value of y so you would make the equations equal to each other like this x²+4x-5=x²-2 then you have add 2 on both side x²+4x-3=x² make it equal to 0 by subtracting x² on both sides 4x-3=0 add 3 on both sides 4x=3 x=3/4

OpenStudy (anonymous):

Im lost where you add 2 on both sides?

OpenStudy (amorfide):

you are moving everything onto one side and making it equal to zero so you do the opposite if you got x²-2 you will subtract x² and add 2 onto both sides

OpenStudy (anonymous):

A. (–3, –8) and (1, 0) B. (–3, –8) and (–1, –4) C. (–3, 0) and (1, –8) D. no solution